# Typical thermalization of low-entanglement states

### v1

@@ Abstract
Proving thermalization from the unitary evolution of a closed quantum system is one of the oldest questions that is still nowadays only partially resolved. Several efforts have led to various formulations of what is called the eigenstate thermalization hypothesis. The latter, however, assume initial states which are highly concentrated around a specific energy window and, as such, cannot account for a large class of states that are of paramount importance and that are operationally accessible in natural physical settings, including many experimental schemes for testing thermalization and for quantum simulation: low-entanglement states. In this work, we prove thermalization of these states under precise conditions that have operational significance. More specifically, we define a random energy smoothing – motivated by arguments of unavoidable finite resolution – on local Hamiltonians that lead to local thermalization when the initial state has low entanglement. Finally we show that such transformation affects neither the Gibbs state locally nor, under a mild condition, the short time dynamics.
@@
Not long after their formulation, it became clear that the postulates of quantum mechanics were to a degree at odds with the principles of statistical mechanics [1]. On the one hand, closed quantum systems evolve unitarily, thus preserving information about their initial conditions. On the other hand, statistical mechanics is centered around concepts such as irreverisibility and thermalization, which allow to describe many-body systems in terms of just a few macroscopic quantities and parameters that depend neither on the initial conditions nor on the microscopic details  [2, 3, 4, 5, 6]. Both theories have nowadays provided correct predictions of uncountable experimental observations. Thermalization of closed many-body quantum systems is in particular observed to occur in practice with overwhelming evidence, also in quantum simulators which allow to probe their dynamics with high levels of precision [7, 8, 9, 10]. This apparent contradiction is eased once one only requires the expectation value of a subset of “physical” observables, for example local observables, to agree with the predictions of statistical mechanics. As a matter of fact, in a lattice system, the state of the system reduced to local patches can relax to a thermal state while the global state evolves unitarily and remains pure. Nevertheless, a derivation of thermalization in this sense from the microscopic description of the dynamics has remained largely elusive. Efforts in this directions have led to various formulations of the eigenstate thermalization hypothesis (ETH) [11, 12], which posits that each eigenstate of a thermalizing Hamiltonian provides the same expectation values of physical observables as one provided by local thermal states. While the ETH is undoubtedly an important milestone, the fact that it requires very precise conditions and assumptions still implies that it cannot account for the vast majority of physical systems where thermalization is observed. Absence of thermalization, as observed in integrable or localized systems [13, 14, 15], seems in fact to be the one requiring particular, carefully set-up conditions. This has led to the idea of typicality, i.e., that “most” physical systems obey statistical mechanics. This usually means that if the system (for instance, its Hamiltonian) is drawn at random from a reasonable set, with overwhelming probability thermalization occurs as expected, and while any given system under examination is usually not randomly drawn, it should behave like a typical one unless explicitly engineered not to. Works in this direction [16, 17, 18, 19, 20] have contributed to formalize the idea that thermalization is the rule, rather than the exception, even in isolated systems.
@@
A drawback of many of these approaches, in the case of both ETH and typicality, is that they often require the assumption that the system is prepared in an initial state which is concentrated around a given energy $E$, i.e., that it has vanishing overlap with energy eigenstates outside of a narrow micro-canonical window centered at $E$. Such states generally cannot be efficiently prepared and, crucially, they exclude important classes of physical and operationally meaningful initial states, such as product states and states with low entanglement, which in addition to being fundamental to the study of lattice systems, are commonly the only accessible initial states to thermalization experiments in quantum simulators or numerical examination of thermalization, for example in the context of probing many-body localization [21, 22, 23, 24, 25].
In this work, we provide this important missing piece of the puzzle by proving thermalization of low-entanglement states. We address this by studying states in an analog of the micro-canonical ensemble which is allowed to have support on the whole energy spectrum. Building upon and extending the work of Ref. [26], we show that, under precise and physically meaningful conditions, these states are locally equivalent to Gibbs state. We later show, generalizing the seminal work of Ref. [19], that such states crucially arise from the long time dynamics of low-entanglement initial state under typical Hamiltonians.
@@
More specifically, given a local Hamiltonian $H$ and an initial low-entanglement state $\rho$, we define an ensemble of random Hamiltonians such that, if $H^{\prime}$ is a Hamiltonian drawn from this ensemble: 1. the Gibbs states of $H$ and $H^{\prime}$ are locally indistinguishable, and 2. with overwhelming probability the equilibrium state resulting from the unitary evolution $e^{iH^{\prime}t}\rho e^{-iH^{\prime}t}$ is locally indistinguishable from said Gibbs state. In addition, we give an assumption under which the dynamics of $H$ and $H^{\prime}$ are indistinguishable in the short time regime. This new kind of Hamiltonian-typicality approach is rooted in physical grounds: it is unrealistic to have a system specified to infinite precision. This approach relaxes and overcomes the common idea that thermalization has to be proven for all Hamiltonians, but rather, it turns out to be typical for an overwhelming portion of meaningful settings.
@@
Throughout this work we consider a cubic $D$-dimensional lattice $\Lambda$ with $N$ sites. With each site $i$ we associate a Hilbert space $\mathcal{H}_{i}$ of dimension $d$. $H$ is a $k$-local Hamiltonian on $\Lambda$ for a constant $k$, meaning that it is of the form $H=\sum_{i\in\Lambda}h_{i}$ where $h_{i}$ has operator norm $1$ and is only supported on sites $j$ s.t. $\mathrm{d}(i,j)\leq k$. Here, $\mathrm{d}(\cdot,\cdot)$ is the standard Manhattan distance on the lattice. We will denote the spectral decomposition of $H$ as $H=\sum_{\nu}E_{\nu}P_{\nu}$. For ease of presentation, in the main text we will assume the eigenstates to be non-degenerate, i.e., $\mathrm{tr}\left(P_{\nu}\right)=1$ for all $E_{\nu}$. In the Supplemental Material, we prove our results relaxing this assumptions. Before discussing thermalization, we need to discuss a natural prerequisite: equilibration [6, 27, 28, 29]. Intuitively, a state equilibrates if after a finite amount of time the system reaches an invariant state. This means that if one looks at the whole evolution as time tends to infinity, the system will spend most of its time close to equilibrium, and hence the state at equilibrium should agree with the infinite-time averaged state
$\rho_{\infty}^{H}=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}\mathrm{d}t\,\rho(t),$ (1)
with $\rho(t)=e^{iHt}\rho e^{-iHt}$. The system is then said to equilibrate if the fluctuations around the average for an observable $A$ are small in the sense that
$\Delta A_{\infty}:=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}\mathrm{d}t\,\mathrm{tr}\left(A(\rho(t)-\rho_{\infty}^{H})\right)^{2}\xrightarrow{N\to\infty}0.$ (2)
@@
$g_{\beta}(H)=\frac{e^{-\beta H}}{Z},\quad Z=\mathrm{tr}\left(e^{-\beta H}\right).$ (3)
In particular, we will be interested in understanding how distinguishable a given state is from a Gibbs state, given only access to local observables. We then introduce the quantity
@@
where $\mathcal{C}_{l}$ denotes the set of all hypercubes in the lattice of side length $l$. If $\rho,\sigma$ are translationally invariant, this is simply equal to the trace norm $\|\rho_{C}-\sigma_{C}\|_{1}$ for any $C\in\mathcal{C}_{l}$, otherwise it measures the distinguishability of $\rho$ and $\sigma$ given access to observables of the form $\frac{1}{|\mathcal{C}_{l}|}\sum_{C\in\mathcal{C}_{l}}O_{C}$, such as typical average local quantities in statistical physics. We will then say that a state is locally thermal if $D_{l}(\rho,g_{\beta}(H))$ converges to 00 with $N$ for some $\beta$. In the remainder of this work, we will make some assumptions on the Gibbs state of the system under examination, $g_{\beta}(H)$. First, we will require it to have exponential decay of correlations. A state $\rho$ is said to have exponential decay of correlations if for some $\xi>0$, for any two regions $X,Y\subset\Lambda$ and $A,B$ supported on $X$ and $Y$ respectively,
$|\langle AB\rangle_{\rho}-\langle A\rangle_{\rho}\langle B\rangle_{\rho}|\leq\|A\|\|B\|e^{-\mathrm{d}(X,Y)/\xi}.$ (5)
The Gibbs state $g_{\beta}(H)$ always has a exponential decay of correlations in one dimension [32], as well as in any higher dimension above a certain critical temperature[33] that depends only on a few parameters of the system. We will denote by $\sigma$ the standard deviation of the energy, i.e., $\sigma^{2}=\mathrm{tr}\left(g_{\beta}(H)H^{2}\right)-\mathrm{tr}\left(g_{\beta}(H)H\right)^{2}$, and assume $\sigma^{2}\geq\Omega(N)$, which implies that the specific heat capacity is non-zero in the thermodynamic limit.
@@
$\omega:=\omega(E,\Delta,\mbox{\boldmath$\delta$},\mathbf{p})=\sum_{k=1}^{K}p_{k}\omega_{\delta_{k}}$ (6)
where $\omega_{\delta_{k}}$ is the micro-canonical ensemble supported inside the window $k$, and where $\mathbf{p}=(p_{1},\ldots,p_{K})$ such that $\sum_{k=1}^{K}p_{k}=1$.
This state therefore physically represents a statistical combination of micro-canonical ensembles; see Fig. 1. For the sake of simplicity, we will choose $\delta_{k}=\delta$ from now on. In the Supplemental Material we show, however, that all our results still hold true if this assumption is relaxed. Before stating our first main result, we need to introduce the notion of the Berry-Esseen (BE) error, which quantifies the difference between a state written in the energy eigenbasis and a Gaussian distribution. More specifically, if $\Pi_{x}$ is the projector onto all energy eigenstates with energy smaller than $x$, then the BE error of $\rho$ with respect to $H$ is defined as $\zeta_{N}=\sup_{x}|\mathrm{tr}\left(\rho\Pi_{x}\right)-G(x)|$, where $G(x)$ is the Gaussian distribution with the same mean and variance as $\rho$. It was proven that if $\rho=g_{\beta}(H)$ for a local Hamiltonian $H$, then $\zeta_{N}\leq O(N^{-1/2})$ [34]. Simple examples saturating this bound are known, and this bound is expected to be saturated by certain non-thermalizing models; under some more generic constraints, however, such as highly entangled eigenstates, a more favorable scaling is expected, even up to $\zeta_{N}\leq e^{-\Omega(N)}$ [35]. From now on we denote by $\zeta_{N}$ the BE error with $\rho=g_{\beta}(H)$. In what follows, we will assume $\zeta_{N}\leq O(N^{-1/2-\kappa})$ for some $\kappa\geq 0$. This includes the standard case $\kappa=0$.
@@
Let $H$ be a local Hamiltonian and $\beta$ be an inverse temperature for which the Gibbs state $g_{\beta}(H)$ has exponential decay of correlations, standard deviation $\sigma\geq\Omega(\sqrt{N})$ and Berry-Esseen error $\zeta_{N}\leq O(N^{-1/2-\kappa})$ for $\kappa\geq 0$. Let $\omega$ denote a GmE with $\Delta,\delta$ satistying
$\mathrm{e}^{\Delta^{2}/\sigma^{2}}\leq N^{\frac{1-\alpha}{D+1}},\quad N^{\frac{1-\alpha}{D+1}-\kappa}\leq\delta\leq\sigma,$ (7)
@@
This first main result shows that, for appropriate choices $\Delta$ and $\delta$, GmE states are locally indistinguishable from Gibbs states. A GmE state can be seen as a mixture of micro-canonical states spanning a range of temperatures, Theorem 1 shows that as long as the range is small enough, the state still looks thermal with a well defined temperature. Notice that if $\kappa>0$, i.e., if the BE error is better than the worst case scenario, $\delta$ can be chosen to decay with the system size.
Keeping in mind our initial goal of capturing equilibrium states resulting from natural and physically motivated initial states, it may seem artificial to consider block-like states with sharp jumps between energy intervals. This is however not the case, as we demonstrate in detail in the remaining of our work. Postponing the discussion about their physicality to the next Section below, we first of all prove that the same ensemble equivalence holds if the state’s structure gets more relaxed, i.e., if it is only approximately GmE in the sense precisely elucidated below.
@@
Let $E,\Delta,\mbox{\boldmath$\delta$}$ as in Definition 1. We define $\omega_{\epsilon}$ an approximate GmE if it is of the form
$\omega_{\epsilon}=p_{\Delta}\left(\sum_{k=1}^{K}p_{k}\tilde{\omega}_{\delta_{k}}\right)+(1-p_{\Delta})\rho_{\mathrm{tail}}$ (9)
@@
$\sum_{k=1}^{K}p_{k}(S(\omega_{\delta_{k}})-S(\tilde{\omega}_{\delta_{k}}))\leq\epsilon,$ (10)
@@
This state represents a more physical version of a GmE state inside the energy window $\Delta$, with decaying tails outside, that has an entropy $\epsilon$ close to the maximum one. Importantly, in the Supplemental Material we demonstrate that Theorem 1 holds true also for the approximate GmE and takes the form
$D_{l}(\omega_{\epsilon},g_{\beta}(H))\leq C_{2}N^{-\gamma_{2}\alpha}+(1-p_{\Delta}),$ (11)
with $\epsilon\leq N^{\frac{1-\alpha}{D+1}}$.
@@
Thermalisation via energy smoothing.
The next question is whether (approximate) GmE states can actually be obtained from Hamiltonian evolution of isolated systems under natural, or typical, conditions. There are two main aspects we consider when talking about “natural” conditions: (i) the Hamiltonian responsible for the time evolution; (ii) the initial state of the system. Regarding (i), we consider typical Hamiltonians in a sense that we will make precise below, in order to exclude edge cases or fine-tuned Hamiltonians for which one does not expect thermalization (for instance, integrable models). Concerning (ii), previous typicality approaches have assumed the initial state to be confined in a well defined energy interval, and have shown properties of the relaxation towards a micro-canonical ensemble in said interval. Here, instead, we start from the assumption of exponential decay of correlations, i.e., low entanglement between spatially separated regions, which we take as natural starting states for lattice systems. These states have been shown to have fast decaying tails in energy [36] which makes them ideal candidates to flow to approximately GmE states. Here, for simplicity of presentation, we focus on the case of product states, and leave the more general case of states with exponential decay of correlations and the proofs to the Supplemental Material, Section II.
@@
Let $\rho$ be a product state and $H$ be a local Hamiltonian. Let $\rho_{\infty}^{UHU^{\dagger}}$ defined as in Eq. 1, where $U$ is drawn from $\mathcal{E}(\delta)$. Consider the interval $I=[E-\Delta,E+\Delta]$ around $E=\mathrm{tr}\left(\rho UHU^{\dagger}\right)$ with $\Delta\geq\omega(\sqrt{N})$ an integer multiple of $\delta$, then
$\rho_{\infty}^{UHU^{\dagger}}=p_{\Delta}\left(\sum_{k:I_{k}\subset I}p_{k}\tilde{\omega}_{\delta_{k}}\right)+(1-p_{\Delta})\rho_{\mathrm{tail}}$ (12)
@@
$\sum_{k:I_{k}\subset I}p_{k}(S(\omega_{\delta_{k}})-S(\tilde{\omega}_{\delta_{k}}))\leq r,$ (13)
where $c_{1},c_{2}$ are system-size independent constants.
@@
Let $H$ be a $k$-local Hamiltonian and $\rho$ be a product state. Let $g_{\beta}(H)$ be the Gibbs state of $H$ at inverse temperature $\beta$ such that $|\mathrm{tr}\left(g_{\beta}(H)H\right)-\mathrm{tr}\left(\rho H\right)|\leq\sigma$. Assume $g_{\beta}(H)$ has exponential decay of correlations, $\sigma\geq\Omega(N)$, and $\zeta_{N}\leq O(N^{-1/2-\kappa})$. For any constant $\alpha\in[0,1)$, choosing $\delta=N^{\frac{1-\alpha}{D+1}-\kappa}$, then with probability at least $1-\exp(-c_{2}N^{\frac{1-\alpha}{D+1}})$ drawing $U$ at random from $\mathcal{E}(\delta)$, we have
$D_{l}(\rho_{\infty}^{UHU^{\dagger}},g_{\beta}(H))\leq C_{2}\,N^{-\gamma_{2}\alpha}+N^{-\gamma_{3}(1-\alpha)},$ (14)

### v2

@@ Abstract
Proving thermalization from the unitary evolution of a closed quantum system is one of the oldest questions that is still nowadays only partially resolved. Several efforts have led to various formulations of what is called the eigenstate thermalization hypothesis, which leads to thermalization under certain conditions on the initial states. These conditions, however, are sensitive to the precise formulation of the hypothesis. In this work, we focus on the important case of low entanglement initial states, which are operationally accessible in many natural physical settings, including experimental schemes for testing thermalization and for quantum simulation. We prove thermalization of these states under precise conditions that have operational significance. More specifically, motivated by arguments of unavoidable finite resolution, we define a random energy smoothing on local Hamiltonians that leads to local thermalization when the initial state has low entanglement. Finally we show that such a transformation affects neither the Gibbs state locally nor, under generic smoothness conditions on the spectrum, the short-time dynamics.
@@
Not long after their formulation, it became clear that the postulates of quantum mechanics were to a degree at odds with the principles of statistical mechanics [1]. On the one hand, closed quantum systems evolve unitarily, thus preserving information about their initial conditions. On the other hand, statistical mechanics is centered around concepts such as irreversibility and thermalization, which allow to describe many-body systems in terms of just a few macroscopic quantities and parameters that depend neither on the initial conditions nor on the microscopic details  [2, 3, 4, 5, 6]. Both theories have nowadays provided correct predictions of uncountable experimental observations. Thermalization of closed many-body quantum systems is in particular observed to occur in practice with overwhelming evidence, also in quantum simulators which allow to probe their dynamics with high levels of precision [7, 8, 9, 10, 11, 12]. This apparent contradiction is eased once one only requires the expectation value of a subset of “physical” observables, for example local observables, to agree with the predictions of statistical mechanics. As a matter of fact, in a lattice system, the state of the system reduced to local patches can relax to a thermal state while the global state evolves unitarily and remains pure. Nevertheless, a derivation of thermalization in this sense from the microscopic description of the dynamics has remained largely elusive. Efforts in this directions have led to various formulations of the eigenstate thermalization hypothesis (ETH) [13, 14], which posits that each eigenstate of a thermalizing Hamiltonian provides the same expectation values of physical observables as the ones provided by local thermal states. The ETH, however, is difficult to verify starting from a microscopic model. In addition, the precise conditions on the initial state and on the observables leading to thermalization are highly sensitive to the formulation of the hypothesis. On the other hand, absence of thermalization, as observed in integrable or localized systems [15, 16, 17], seems to be the occurrence requiring particular, carefully set-up conditions. This has led to the idea of typicality, i.e., that “most” physical systems obey statistical mechanics. This usually means that if the system (for instance, its Hamiltonian) is drawn at random from a reasonable set, with overwhelming probability thermalization occurs as expected, and while any given system under examination is usually not randomly drawn, it should behave like a typical one unless explicitly engineered not to. Works in this direction [18, 19, 20, 21, 22] have contributed to formalize the idea that thermalization is the rule, rather than the exception, even in isolated systems.
@@
A drawback of many of these approaches, in the case of both the ETH and typicality, is that they often require the assumption that the system is prepared in an initial state which is concentrated around a given energy $E$. While all states with sub-extensive energy variance are concentrated around their average energy, it is often not clear how strong the concentration needs to be in order for thermalization to be guaranteed given either a formulation of ETH or a typicality transformation. One often assumes that the support of the state vanishes outside of a narrow micro-canonical energy window; depending on how narrow this window is such a state generally cannot be efficiently prepared. Crucially, this excludes important classes of physical and operationally meaningful initial states, such as product states and states with low entanglement. These states, in addition to being fundamental to the study of lattice systems, are commonly the only accessible initial states to thermalization experiments in quantum simulators or numerical examination of thermalization, for example in the context of probing many-body localization [23, 24, 25, 26, 27], and are often considered in works on relaxation and equilibration [28, 29, 30].
@@
In this work, we provide this important missing piece of the puzzle by proving thermalization of low-entanglement states. We address this by studying states in an analog of the micro-canonical ensemble which is allowed to have support on the whole energy spectrum. Building upon and extending the work of Ref. [31], we show that, under precise and physically meaningful conditions, these states are locally equivalent to Gibbs states. We later show, generalizing the seminal work of Ref. [21], that such states crucially arise from the long-time dynamics of low-entanglement initial states under typical Hamiltonians. Our typicality transformation consists of a unitary operation that randomly mixes eigenstates with nearby energies. Related transformations have been introduced in Ref. [32] to compute higher order correlation functions under a generalized version of the ETH stemming from local (in energy) unitary invariance, and this has been connected to the theory of free probability in Ref. [33]. In our work, we explicitly demonstrate how such a transformation leads to equilibrium states being locally indistinguishable from thermal states. Moreover, by bounding the energy difference between eigenstates that get mixed, we elucidate how the original Hamiltonian is related to the transformed one. More specifically, given a local Hamiltonian $H$ and an initial low-entanglement state $\rho$, we define an ensemble of random Hamiltonians such that, if $H^{\prime}$ is a Hamiltonian drawn from this ensemble: 1. the Gibbs states of $H$ and $H^{\prime}$ are locally indistinguishable, and 2. with overwhelming probability the equilibrium state resulting from the unitary evolution $\mathrm{e}^{-\mathrm{i}H^{\prime}t}\rho\mathrm{e}^{\mathrm{i}H^{\prime}t}$ is locally indistinguishable from said Gibbs state. In addition, we give an assumption under which the dynamics of $H$ and $H^{\prime}$ are indistinguishable in the short-time regime. This new kind of Hamiltonian-typicality approach is rooted in physical grounds: it is unrealistic to have a system specified to infinite precision. This approach relaxes and overcomes the common idea that thermalization has to be proven for all Hamiltonians, but rather, it turns out to be typical for an overwhelming portion of meaningful settings.
@@
Throughout this work we consider a cubic $D$-dimensional lattice $\Lambda$ with $N$ sites. With each site $i$ we associate a Hilbert space $\mathcal{H}_{i}$ of dimension $d$. $H$ is a $k$-local Hamiltonian on $\Lambda$ for a constant $k$, meaning that it is of the form $H=\sum_{i\in\Lambda}h_{i}$ where $h_{i}$ has operator norm $1$ and is only supported on sites $j$ s.t. $\mathrm{d}(i,j)\leq k$. Here, $\mathrm{d}(\cdot,\cdot)$ is the standard Manhattan distance on the lattice. We will denote the spectral decomposition of $H$ as $H=\sum_{\nu}E_{\nu}P_{\nu}$. For ease of presentation, in the main text we will assume the eigenstates to be non-degenerate, i.e., $\mathrm{tr}\left(P_{\nu}\right)=1$ for all $E_{\nu}$. In the Supplemental Material, we prove our results relaxing this assumptions. We will use standard $O$ and $\Omega$ notation, and we will use $\tilde{O},\tilde{\Omega}$ to denote asymptotic upper bounds where logarithmic factors are ignored, i.e. $f(h)=\tilde{O}(g(h))$ iff $f(h)=O(\log^{r}(h)g(h))$ for some $r$. Before discussing thermalization, we need to discuss a natural prerequisite: equilibration [6, 34, 35, 36]. Intuitively, a state equilibrates if after a finite amount of time the system reaches an invariant state. This means that if one looks at the whole evolution as time tends to infinity, the system will spend most of its time close to equilibrium, and hence the state at equilibrium should agree with the infinite-time averaged state
$\rho_{\infty}^{H}=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}\mathrm{d}t\,\rho^{H}(t),$ (1)
with $\rho^{H}(t)=\mathrm{e}^{iHt}\rho\mathrm{e}^{-iHt}$. The system is then said to equilibrate if the fluctuations around the average for an observable $A$ are small in the sense of
$\Delta A_{\infty}:=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}\mathrm{d}t\,\mathrm{tr}\left(A(\rho^{H}(t)-\rho_{\infty}^{H})\right)^{2}\xrightarrow{N\to\infty}0.$ (2)
@@
$g_{\beta}(H)=\frac{\mathrm{e}^{-\beta H}}{Z},\quad Z=\mathrm{tr}\left(\mathrm{e}^{-\beta H}\right).$ (3)
In particular, we will be interested in understanding how distinguishable a given state is from a Gibbs state, given only access to local observables. We introduce the quantity
@@
where $\mathcal{C}_{l}$ denotes the set of all hypercubes in the lattice of side length $l$ and $\rho_{C}$ is the state $\rho$ reduced to $C$. If $\rho,\sigma$ are translationally invariant, this is simply equal to the trace norm $\|\rho_{C}-\sigma_{C}\|_{1}$ for any $C\in\mathcal{C}_{l}$. Otherwise it measures the distinguishability of $\rho$ and $\sigma$ given access to observables of the form $\frac{1}{|\mathcal{C}_{l}|}\sum_{C\in\mathcal{C}_{l}}O_{C}$, such as typical average local quantities in statistical physics. We will then say that a state is locally thermal if $D_{l}(\rho,g_{\beta}(H))$ converges to 00 with $N$ for some $\beta$. In the remainder of this work, we will make some assumptions on the Gibbs state of the system under examination, $g_{\beta}(H)$. First, we will require it to have exponential decay of correlations. A state $\rho$ is said to have exponential decay of correlations if for some $\xi>0$, for any two regions $X,Y\subset\Lambda$ and $A,B$ supported on $X$ and $Y$ respectively,
$|\langle AB\rangle_{\rho}-\langle A\rangle_{\rho}\langle B\rangle_{\rho}|\leq\|A\|\|B\|\mathrm{e}^{-\mathrm{d}(X,Y)/\xi}.$ (5)
The Gibbs state $g_{\beta}(H)$ always has an exponential decay of correlations in one dimension [39], as well as in any higher dimension above a certain critical temperature [40] that depends only on a few parameters of the system. We will denote by $\sigma$ the standard deviation of the energy, i.e., $\sigma^{2}=\mathrm{tr}\left(g_{\beta}(H)H^{2}\right)-\mathrm{tr}\left(g_{\beta}(H)H\right)^{2}$, and assume $\sigma^{2}\geq\Omega(N)$, which implies that the specific heat capacity is non-zero in the thermodynamic limit.
@@
$\omega:=\omega(E,\Delta,\mbox{\boldmath$\delta$},\mathbf{q})=\sum_{k=1}^{K}q_{k}\omega_{\delta_{k}}$ (6)
where $\omega_{\delta_{k}}$ is the micro-canonical ensemble supported inside the window $k$, and where $\mathbf{q}=(q_{1},\ldots,q_{K})$ such that $\sum_{k=1}^{K}q_{k}=1$.
This state therefore physically represents a statistical combination of micro-canonical ensembles; see Fig. 1. For the sake of simplicity, we will choose $\delta_{k}=\delta$ from now on. However, in the Supplemental Material we show that all our results still hold true if this assumption is relaxed. Before stating our first main result, we need to introduce the notion of the Berry-Esseen (BE) error, which quantifies the difference between a state written in the energy eigenbasis and a Gaussian distribution. More specifically, if $\Pi_{x}$ is the projector onto all energy eigenstates with energy smaller than $x$, then the BE error of $\rho$ with respect to $H$ is defined as $\zeta_{N}=\sup_{x}|\mathrm{tr}\left(\rho\Pi_{x}\right)-G(x)|$, where $G(x)$ is the Gaussian distribution with the same mean and variance as $\rho$. It was proven that if $\rho$ has exponential decay of correlations, then $\zeta_{N}\leq\tilde{O}(N^{-1/2})$ [31]. Simple examples saturating this bound are known, and this bound is expected to be saturated by certain non-thermalizing models. Nonetheless, under some more generic constraints, such as highly entangled eigenstates, a more favorable scaling is expected, even up to $\zeta_{N}\leq\mathrm{e}^{-\Omega(N)}$ [41]. From now on we denote by $\zeta_{N}$ the BE error with $\rho=g_{\beta}(H)$. In what follows, we will assume $\zeta_{N}\leq\tilde{O}(N^{-1/2-\kappa})$ for some $\kappa\geq 0$. This includes the worst case $\kappa=0$.
@@
Let $H$ be a local Hamiltonian and $\beta$ an inverse temperature for which the Gibbs state $g_{\beta}(H)$ has exponential decay of correlations, standard deviation $\sigma\geq\Omega(\sqrt{N})$ and Berry-Esseen error $\zeta_{N}\leq\tilde{O}(N^{-1/2-\kappa})$ for $\kappa\geq 0$. Let $\omega$ denote a GmE with $\Delta,\delta$ satisfying
$\mathrm{e}^{\Delta^{2}/\sigma^{2}}\leq\tilde{O}\left(N^{\frac{1-\alpha}{D+1}}\right),\quad\Omega\left(N^{\frac{1-\alpha}{D+1}-\kappa}\right)\leq\delta\leq\sigma,$ (7)
@@
This first main result shows that, for appropriate choices $\Delta$ and $\delta$, GmE states are locally indistinguishable from Gibbs states. A GmE state can be seen as a mixture of micro-canonical states spanning a range of temperatures and Theorem 1 shows that as long as its range is small enough, the state still looks thermal with a well-defined temperature. Notice that if $\kappa>0$, i.e., if the BE error is better than the worst case scenario, $\delta$ can be chosen to decay with the system size.
Keeping in mind our initial goal of capturing equilibrium states resulting from natural and physically motivated initial states, it may seem artificial to consider only block-like states with sharp jumps between energy intervals. Therefore, postponing the discussion about their physicality to the next Section below, we first of all prove that the same ensemble equivalence holds if the state’s structure gets more relaxed, i.e., if it is only approximately GmE in the sense precisely elucidated below.
@@
Let $E,\Delta,\mbox{\boldmath$\delta$},\mathbf{q}$ be as in Definition 1. We define $\omega_{\eta}$ an approximate GmE if it is of the form
$\omega_{\eta}=p_{\Delta}\left(\sum_{k=1}^{K}q_{k}\tilde{\omega}_{\delta_{k}}\right)+(1-p_{\Delta})\rho_{\mathrm{tail}}$ (9)
@@
$\sum_{k=1}^{K}q_{k}(S(\omega_{\delta_{k}})-S(\tilde{\omega}_{\delta_{k}}))\leq\eta,$ (10)
@@
This state represents a more physical version of a GmE state inside the energy window $\Delta$, with decaying tails outside, that has an entropy $\eta$-close to the maximum one. Importantly, in the Supplemental Material we demonstrate that Theorem 1 holds true also for the approximate GmE and takes the form
$D_{l}(\omega_{\eta},g_{\beta}(H))\leq C_{2}N^{-\gamma_{2}\alpha}+2(1-p_{\Delta}),$ (11)
with $\eta\leq N^{\frac{1-\alpha}{D+1}}$.
@@
Thermalization via energy smoothing.
The next question is whether (approximate) GmE states can actually be obtained from Hamiltonian evolution of isolated systems under natural, or typical, conditions. There are two main aspects we consider when talking about “natural” conditions: (i) the Hamiltonian responsible for the time evolution; (ii) the initial state of the system. Regarding (i), we consider typical Hamiltonians in a sense that we will make precise below, in order to exclude edge cases or fine-tuned Hamiltonians for which one does not expect thermalization (for instance, integrable models). Concerning (ii), previous typicality approaches have assumed the initial state to be confined in a well-defined energy interval, and have shown properties of the relaxation towards a micro-canonical ensemble in said interval. Here, instead, we start from the assumption of exponential decay of correlations, i.e., low entanglement between spatially separated regions, which we take as natural starting states for lattice systems. These states have been shown to have fast decaying tails in energy [42] which makes them ideal candidates to flow to approximately GmE states. Here, for simplicity of presentation, we focus on the case of product states, and leave the more general case of states with exponential decay of correlations and the proofs to the Supplemental Material, Section II.
@@
Let $\rho$ be a product state and $H$ be a local Hamiltonian. Let $\rho_{\infty}^{UHU^{\dagger}}$ be defined as in Eq. 1, where $U$ is drawn from $\mathcal{E}(\delta)$. Consider the interval $I=[E-\Delta,E+\Delta]$ around $E=\mathrm{tr}\left(\rho UHU^{\dagger}\right)$ with $\Delta\geq\omega(\sqrt{N})$ an integer multiple of $\delta$, then
$\rho_{\infty}^{UHU^{\dagger}}=p_{\Delta}\left(\sum_{k:I_{k}\subset I}q_{k}\tilde{\omega}_{\delta_{k}}\right)+(1-p_{\Delta})\rho_{\mathrm{tail}}$ (12)
@@
$\sum_{k:I_{k}\subset I}q_{k}(S(\omega_{\delta_{k}})-S(\tilde{\omega}_{\delta_{k}}))\leq r,$ (13)
where $c_{1}$ is a system-size independent constant.
@@
Let $H$ be a $k$-local Hamiltonian and $\rho$ be a product state. Let $g_{\beta}(H)$ be the Gibbs state of $H$ at inverse temperature $\beta$ such that $|\mathrm{tr}\left(g_{\beta}(H)H\right)-\mathrm{tr}\left(\rho H\right)|\leq\sigma$. Assume $g_{\beta}(H)$ has exponential decay of correlations, $\sigma\geq\Omega(\sqrt{N})$, and $\zeta_{N}\leq\tilde{O}(N^{-1/2-\kappa})$. For any constant $\alpha\in[0,1)$, choosing $\delta=\Omega\left(N^{\frac{1-\alpha}{D+1}-\kappa}\right)$, with probability at least $1-\exp(-c_{2}N^{\frac{1-\alpha}{D+1}})$ drawing $U$ at random from $\mathcal{E}(\delta)$, we have
$D_{l}(\rho_{\infty}^{UHU^{\dagger}},g_{\beta}(H))\leq C_{2}\,N^{-\gamma_{2}\alpha}+\tilde{O}\left(N^{-\gamma_{3}(1-\alpha)}\right),$ (14)